OpenStudy (anonymous):

Write each fraction in simplest form. 8/24

OpenStudy (bibby):

factors of 8 1 2 4 8 factors of 24 1 2 3 4 8 6 12 24 Use the GCF

OpenStudy (anonymous):

use a calculator or just do 24/8 = 3 so if you have 8/24 it's written as 1/3

OpenStudy (anonymous):

Learn to make factor trees 8=2*2*2 24=2*2*2*3

OpenStudy (anonymous):

Thank you @BizPro @Rouault @bibby